OpenStudy (shadowcain):

Solve for x. 2x - 4 2x --------- = ----- 10 20

OpenStudy (denonakavro):

Do u know how to cross multiply

OpenStudy (shadowcain):

yes

OpenStudy (denonakavro):

So multiply 2x-4 with 20 and that would equal 10 times 2x

OpenStudy (shadowcain):

oh that really helped

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \frac{2x-4 }{10} =\frac{2x}{20} }\) multiply times 2 on top and bottom \(\large\color{#000000 }{ \displaystyle \frac{4x-8 }{20} =\frac{2x}{20} }\) And therefore; \(\large\color{#000000 }{ \displaystyle 4x-8 =2x }\) (if you want to get creative)

OpenStudy (solomonzelman):

same way, you can divide by 2 on top and bottom on the right side (instead of multiplying times 2 on top and bottom on the left side).